Strands of copper wire from a manufacturer are analyzed for strength and conductivity. The results from 100 strands are as follows: th High conductivity 70 17 Low conductivity103 a) If a strand is randomly selected, what is the probability that its conductivity is high and its strength is high? Round your answer to two decimal places (e-g. 98.76). b) If a strand is randomly selected, what is the probability that its conductivity is low or the strength is low? Round your answer to two decimal places (e.g. 98.76). c) Consider the event that a strand has low conductivity and the event that the strand has a low strength. Are these two events mutually exclusive?
Answer #1

a)
P(high conductivity and high strength)
= 70/100
=0.70

b)
P(low conductivity or low strength)
= (10 + 3 +17)/100
=0.30

c)

P(low conductivity and low strength)
= 3/100
which is not 0
hence both are not mutually exclusive